OpenStudy (anonymous):

Researchers studying fish populations in Sunset Lake caught, marked, and then released 232 bass. Later, a sample of 329 bass were caught and examined. Of these, 16 were found to be marked. Use the proportion below to estimate the total bass population in Sunset Lake.

OpenStudy (anonymous):

marked bass in sample/ total bass in sample= marked bass and lake/ total bass in lake

OpenStudy (anonymous):

So set up the proportion as: \[\frac{ mark_{sample} }{ total_{caught} } = \frac{ total_{sampled} }{ all_{fish} } \rightarrow \frac{ 16 }{ 329 } = \frac{ 232 }{ x}\] Ans just solve for x.

OpenStudy (anonymous):

its 4770 right